OpenStudy (anonymous):

What values for theta (0

OpenStudy (anonymous):

@AQ99

OpenStudy (aq99):

hmm...which do you think it is? c:

OpenStudy (anonymous):

Well it's to solve for each of those equations >.< The options for equation a are a. pi/3, 5 pi/3 b. 2pi/3, 4pi/3 c. pi/6, 3pi/6 d. 7pi/6, 11pi/6

OpenStudy (anonymous):

And the options for equation b are a. pi/2, 2pi/2 b. 3pi/4, 5pi/4 c. pi/3, 5pi/3 d. 2pi/3, 4pi/3

OpenStudy (aq99):

>.< oomg

OpenStudy (anonymous):

@iGreen @Hero Please please help!!

OpenStudy (aq99):

@iGreen @Hero

OpenStudy (anonymous):

^^^great minds think alike

OpenStudy (aq99):

XD we both tagged him at the same time! :DD

OpenStudy (anonymous):

@jim_thompson5910 help?

OpenStudy (aq99):

D: @Hero !!

jimthompson5910 (jim_thompson5910):

the equation for part A is this right? \[\Large 3\sin(\theta) = \sin(\theta) - 1\]

OpenStudy (anonymous):

Yessir

jimthompson5910 (jim_thompson5910):

the goal is to isolate theta first isolate sin(theta) how would you isolate sin(theta) ?

jimthompson5910 (jim_thompson5910):

hint: let z = sin(theta) that allows us to make substitutions to get 3z = z-1 what do you get when you solve for z ?

OpenStudy (anonymous):

2z=-1? So z=-1/2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so sin(theta) = -1/2

jimthompson5910 (jim_thompson5910):

what values of theta make the equation \[\Large \sin(\theta) = -\frac{1}{2}\] true?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

here's a unit circle to help answer the question http://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/1024px-Unit_circle_angles_color.svg.png locate any points that have -1/2 as a y coordinate. The corresponding angle will be a solution for theta

OpenStudy (anonymous):

11pi/6?

OpenStudy (anonymous):

and 7pi/6?

OpenStudy (anonymous):

Can you help with equation b too?

jimthompson5910 (jim_thompson5910):

yep 7pi/6 and 11pi/6

jimthompson5910 (jim_thompson5910):

for equation b, you'll use the identity \[\Large \tan^2(\theta) = \sec^2(\theta) - 1\]

OpenStudy (aq99):

Thanks so much for the help, jim. I was at a huge loss here :/ sorry I couldnt help out.

jimthompson5910 (jim_thompson5910):

\[\Large \tan^2(\theta) = -\frac{3}{2}\sec(\theta)\] \[\Large \sec^2(\theta)-1 = -\frac{3}{2}\sec(\theta)\] \[\Large z^2-1 = -\frac{3}{2}z\] \[\Large 2z^2-2 = -3z\] \[\Large 2z^2+3z-2 = 0\] On the third line I replaced every sec(theta) with z. I let z = sec(theta). From here you need to solve for z (use the quadratic formula)

OpenStudy (anonymous):

z=-2? @jim_thompson5910

OpenStudy (anonymous):

sorry that took a second /.\

jimthompson5910 (jim_thompson5910):

you'll also find that z = 1/2, but sec(theta) = 1/2 has no solutions. So we can ignore that part.

jimthompson5910 (jim_thompson5910):

yes z = -2 which leads to sec(theta) = -2 ---> cos(theta) = -1/2

OpenStudy (anonymous):

Okay so what do I do next?

jimthompson5910 (jim_thompson5910):

look for points on the unit circle that have an x coordinate of -1/2

jimthompson5910 (jim_thompson5910):

since points of the unit circle have x coordinates corresponding to cos(theta)

OpenStudy (anonymous):

2pi/3 and 4pi/3?

jimthompson5910 (jim_thompson5910):

correct